# A Behaviour and Disease Model of Testing and Isolation

### v1

@@ Abstract
There has been interest in the interactions between infectious disease dynamics and behaviour for most of the history of mathematical epidemiology. This has included consideration of which mathematical models best capture each phenomenon, as well as their interaction, but typically in a manner that is agnostic to the exact behaviour in question. Here, we investigate interacting behaviour and disease dynamics specifically related to behaviours around testing and isolation. This epidemiological-behavioural interaction is of particular interest as, prospectively, it is well-placed to be informed by real-world data temporally monitoring test results and compliance with testing policy. To carry out our investigation we extend an existing “behaviour and disease” (BaD) model by incorporating the dynamics of symptomatic testing and isolation. We provide a dynamical systems analysis of the ordinary differential equations that define this model, providing theoretical results on its behaviour early in a new outbreak (particularly its basic reproduction number) and endemicity of the system (its steady states and associated stability criteria). We then supplement these findings with a numerical analysis to inform how temporal and cumulative outbreak metrics depend on the model parameter values for epidemic and endemic regimes. As the presented interdisciplinary modelling approach can accommodate further extensions (including, but not limited to, adding testing capacity, decay in behavioural effects and multiple pathogen variants), we hope that our work will encourage further modelling studies integrating specific measured behaviours and disease dynamics that may reduce the health and economic impacts of future epidemics.
@@ 1.1 Background
From early in the history of infectious disease modelling, it was clear that the broad mathematical approaches used to model pathogen transmission could also be used to model social transmission such as the spread of rumours or behaviours, and that there were important specific differences between social and viral spreading, theoretically and empirically [1, 2]. As data availability and quality from outbreaks improved, empirical evidence for the impact of behavioural response on infection spread increased, particularly during the 2004 SARS epidemic [3]. A particularly influential 2009 paper by Funk et al. [4] (with over 800 citations at the time of writing) explicitly coupled mechanistic models of infection and behavioural response to infection. Even in 2010, a systematic review of approaches to this problem had 85 citations in its bibliography [5].
In the last 15 years there has been significant work on the interaction of behaviour and epidemics including theoretical analyses [6, 7], consideration of social network structure and demographics [8], and attempts to elucidate prototypical models combining both mechanisms [9]. A dedicated monograph on behavioural dynamics was published in 2013 [10], and a further review paper in 2015 [11]. Spreading of behaviour continues to inspire mathematical developments, for example development of differential equation systems with cubic and higher-order interaction terms [12], use of multiplex network theory [13], and investigation of paradoxical behaviour [14]. The COVID-19 pandemic, seeing significant impacts from behaviour, also saw attempts to advance this field [15].
Despite this activity, important challenges remain in developing a generalised model of epidemiological and behavioural dynamics, including balancing model complexity, interpretability and capability to validate them with empirical data [16]. And at the same time, few models have attempted to capture specific behaviours, preferring to adopt an intervention-agnostic approach.
@@ 1.2 Outline of this study
In this paper, we define and analyse an extension to the Behaviour and Disease (BaD) model of [17, 18] to the behavioural dynamics specific to (symptomatic) testing and isolation. This is because testing and isolation are known to have both a strong behavioural component and a significant impact on disease transmission [19, 20]. In addition, in terms of data availability it is also possible to monitor both test results and compliance with testing policy over time using community surveys [21].
We structure the remainder of the paper as follows. We initially present our model that incorporates dynamic testing behaviour and infectious disease transmission processes (Section 2). We then conduct theoretical analyses of the model (Section 3) to give insight on the early dynamics and endemicity of the system; specifically, we derive its basic reproduction number $\mathcal{R}_{0}$ (Section 3.1), compute its steady states (Section 3.2) and analyse stability criteria (Section 3.3). Note that derivation of quantities such as final size in the absence of waning immunity are not believed to be easily available for this class of models using existing methods (Röst, personal communication relating to the model published in [22]). Accordingly, to investigate how temporal and cumulative outbreak metrics depend on model parameter values we describe the results of numerical simulations (Section 4). We conclude by discussing the findings and their implications (Section 5).
@@ 2 Behaviour and Disease (BaD) model for symptomatic testing
In this section we introduce our mathematical model, which extends the BaD model of [17] to incorporate dynamics that model symptomatic testing. The key concept behind this model is the distinction between individuals who do not intend to seek testing when symptomatic, referred to as non-behavers (labelled $N$) and those who seek testing when symptomatic, referred to as behavers (labelled $B$). We emphasise, however, that while we use the nomenclature of behavers and non-behavers for consistency with the other literature, we are considering $B$ individuals as those who hold an intention to test if symptomatic, and so may fail to do this if, for example, there is a lack of testing capacity. This means we parameterise the model in terms of the effectiveness of testing as an intervention, including the real-world factors that can obstruct access to testing, rather than its efficacy under ideal conditions that is primarily determined by test sensitivity.
@@
Our pathogen dynamical framework is built upon the standard SEIRS compartmental framework: susceptible individuals ($S$) become exposed ($E$), progress to an infectious state ($I$, $A$ or $T$ as explained below) and then recover ($R$), with a non-zero rate of waning immunity meaning that recovered individuals can become susceptible again. To account for symptomatic testing behaviour, we further split the infectious class into asymptomatics (or pauci-symptomatics) labelled $A$, and symptomatics who are labelled $I$ if they do not seek or do not receive a positive test, and $T$ if they seek and receive a positive test. The transitions between epidemiological states follow the standard SEIRS framework, while the changes in testing behaviour are governed by the rates of behaviour uptake ($\omega$) and abandonment ($\alpha$). Figure 1 shows the structure of this model.
@@
Let $B(t)$ denote the proportion of the population at time $t$ that would test if showing symptoms, and $N(t)$ denote the proportion that would not. We will henceforth mostly suppress explicit dependence on $t$ for notational compactness. For each state class $X\in\{S,E,I,A,R\}$, $X_{N}$ and $X_{B}$ denote the proportions of non-behavers and behavers, respectively. Additionally, let the proportion of symptomatic individuals who seek testing and test positive be $T$.
@@
We adopt the same functional forms for the rates of behaviour uptake and behaviour loss as the original BaD model [17]. The rate of behaviour uptake is governed by
@@
where $\phi_{1}$ represents the influence of social contagion on testing behaviour, $\phi_{2}$ accounts for behaviour uptake driven by the perception of illness threat, and $\omega_{3}$ captures spontaneous adoption of the behaviour. The rate of behaviour loss is
@@
where $\psi$ represents the influence of social contagion on not seeking testing and $\alpha_{2}$ captures spontaneous abandonment of the behaviour. Note that while, for a complex contagion model, the quadratic and higher-order terms on the right-hand sides of Equations (1) and (2) might be large and even provide the dominant contribution to $\omega$ and $\alpha$, we will only include linear terms in our analysis when an explicit form is required.
@@
whereas for behavers it is $q_{B}\lambda$.
Figure 1: A compartmental diagram illustrating the BaD model for symptomatic testing. We distinguish between individuals who do not seek testing when symptomatic, referred to as non-behavers (labelled $N$) and those who seek testing when symptomatic, referred to as behavers (labelled $B$). Changes in testing behaviour are governed by the rates of behaviour uptake ($\omega$) and abandonment ($\alpha$). The epidemiological states are susceptible ($S$), exposed ($E$), asymptomatic infectious (or pauci-symptomatics) labelled $A$, symptomatic infectious who are labelled $I$ if they do not seek or do not receive a positive test and $T$ if they seek and receive a positive test, and those who are recovered ($R$). The transitions between epidemiological states follow a standard SEIRS framework, with waning immunity where recovered individuals can become susceptible again. We define the model parameters in Table 1.
@@
      $q_{A}$ the reduction in infectiousness from being asymptomatic $0.58$ [24]
      $q_{T}$ the reduction in infectiousness from isolation $0.25$
      $q_{B}$ the reduction in susceptibility from the testing behaviour $0.5$ [17]
@@
$\omega(t)$ the “force of infection" for testing behaviour uptake
@@
$\alpha(t)$ the “force of infection" for testing behaviour abandonment
@@
with $B=1-N$. Linearising about small initial behaviour ($B(0)\ll 1$) we have $N\approx 1$ and, if $\omega_{3}=0$, behaviour spreads through social pressure if $\omega_{1}>\alpha_{1}+\alpha_{2}$. Following [17], we define the social reproduction number to be
@@
$F(N)=\left(\alpha_{1}-\omega_{1}\right)N^{2}-\left(\alpha_{1}-\alpha_{2}-\omega_{1}-\omega_{3}\right)N-\alpha_{2}=0$ (14)
@@ 3.1.1 In the absence of behaviour
If $\omega_{3}=0$ and $\mathcal{R}_{0}^{B}<1$ then the infection-free steady state has $S_{N}=1$ and $S_{B}=0$. The epidemic takes off if
@@ 3.1.2 In the presence of behaviour
We now consider the full model where both behaviour and infection are present (Equations (3)-(13)). Here we calculate the next-generation matrix. The model has two states-at-infection, $E_{N}$ and $E_{B}$. Let $t_{N}(X)$ denote the expected time an individual spends in compartment $X$ following state-at-infection $E_{N}$. Similarly, let $t_{B}(X)$ represent the corresponding time spent in compartment $X$ following state-at-infection $E_{B}$.
Defining $\alpha_{0}=\alpha(N_{0})$ and $\omega_{0}=\omega(B_{0},0)$, we derive the following expressions for the expected time spent in each compartment. For individuals with state-at-infection $E_{N}$, we have
$t_{N}(A_{N})$ $=\frac{p_{A}\left[\alpha_{0}\left(\sigma+\alpha_{0}+\omega_{0}\right)+\gamma\left(\sigma+\alpha_{0}\right)\right]}{\gamma\left(\sigma+\alpha_{0}+\omega_{0}\right)\left(\gamma+\alpha_{0}+\omega_{0}\right)}$ (15)
$t_{N}(I_{N})$ $=\frac{\left(1-p_{A}\right)\left[\alpha_{0}\left(\sigma+\alpha_{0}+\omega_{0}\right)+\gamma\left(\sigma+\alpha_{0}\right)-p_{T}\alpha_{0}\omega_{0}\right]}{\gamma\left(\sigma+\alpha_{0}+\omega_{0}\right)\left(\gamma+\alpha_{0}+\omega_{0}\right)}$
@@
The terms $\Lambda_{N}$ and $\Lambda_{B}$ can be further expressed as
@@
where $\Lambda_{N},\Lambda_{B}$ are given by Equation (19) and (20), $N_{0}$ solves Equation (14) and $B_{0}=1-N_{0}.$ Note that $\Lambda_{N}$ and $\Lambda_{B}$ can be rewritten as
@@
where $\mathcal{R}_{0}^{D}$ is the behaviour-free reproduction number. Writing
$\Delta\mathcal{R}=\frac{\beta(1-p_{A})(1-q_{T})p_{T}}{\gamma}$
shows that
$\mathcal{R}_{0}=\underbrace{N_{0}\mathcal{R}_{0}^{D}}_{\text{(i)}}+\underbrace{q_{B}B_{0}\mathcal{R}_{0}^{D}}_{\text{(ii)}}-\underbrace{(1-p_{N})N_{0}\Delta\mathcal{R}}_{\text{(iii)}}-\underbrace{q_{B}p_{B}B_{0}\Delta\mathcal{R}}_{\text{(iv)}}\,.$ (21)
In the Equation (21): (i) captures the spread of infection from non-isolators (those in $A$ and $I$) to non-behavers (those in $N$); (ii) captures the spread of infection from non-isolators to behavers (those in $B$); (iii) captures the reduction in infections in non-behavers caused by isolation (those in $T$), and; (iv) captures the reduction in infections in behavers caused by isolation. Note that in the limiting case where $B_{0}$ approaches 00, $p_{N}$ approaches $1$ and Equation (21) shows that $\mathcal{R}_{0}$ approaches $\mathcal{R}_{0}^{D}$.

### v2

@@ Abstract
There has been interest in the interactions between infectious disease dynamics and behaviour for most of the history of mathematical epidemiology. This has included consideration of which mathematical models best capture each phenomenon, as well as their interaction, but typically in a manner that is agnostic to the exact behaviour in question. Here, we investigate interacting behaviour and disease dynamics specifically related to decisions around testing and isolation. To carry out our investigation we extend an existing “behaviour and disease” (BaD) model by incorporating the dynamics of symptomatic testing and isolation, including the influence of positive tests on perception of infection risk. We provide a dynamical systems analysis of the ordinary differential equations that define this model, providing theoretical results on its behaviour early in a new outbreak (particularly its basic reproduction number) and endemicity of the system (its steady states and associated stability criteria). We then supplement these findings with a numerical analysis to inform how temporal and cumulative outbreak metrics depend on the model parameter values for epidemic and endemic regimes. We observe novel model outputs such as epidemics that have more observed cases detected through increased testing, but are less objectively severe in terms of total number of infections.
@@
Plain language summary
Throughout the history of infectious disease outbreaks, people have modified their behaviour in response to risk. Such behaviours can include vaccination, seeking treatment, and non-pharmaceutical interventions like hand-washing and social distancing. There have been various attempts to include behavioural response quantitatively in models of epidemics, but these have typically not specified the exact behaviour. In this article, we develop and analyse a joint mathematical model of infection and specific decisions around testing and isolation. This epidemiological-behavioural interaction is of particular interest as, prospectively, it is well-placed to be informed by real-world data temporally monitoring test results and compliance with testing policy. As the presented interdisciplinary modelling approach can accommodate further extensions (including, but not limited to, adding testing capacity, decay in behavioural effects and multiple pathogen variants), we hope that our work will encourage further modelling studies integrating specific measured behaviours and disease dynamics that may reduce the health and economic impacts of future epidemics.
@@ 1.1 Background
From early in the history of infectious disease modelling, it was clear that the broad mathematical approaches used to model pathogen transmission could also be used to model social transmission such as the spread of rumours or behaviours, and that there were important specific differences between social and viral spreading, theoretically and empirically [1, 2]. As data availability and quality from outbreaks improved, empirical evidence for the impact of behavioural response on infection spread increased, particularly during the 2004 SARS epidemic [3].
A prominent 2009 paper by Funk et al. [4] influenced much subsequent research by providing a framework to consider the interaction of viral and behavioural spreading. In the last 15 years there has been significant progress on modelling the interaction of behaviour and epidemics including theoretical analyses [5, 6], consideration of social network structure and demographics [7], and attempts to elucidate prototypical models combining both mechanisms [8]. An early review of the field was published in 2010 [9], a dedicated monograph on behavioural dynamics was published in 2013 [10], and a further review paper in 2015 [11]. Spreading of behaviour continues to inspire mathematical developments, for example development of differential equation systems with cubic and higher-order interaction terms [12], use of multiplex network theory [13], and investigation of paradoxical behaviour [14]. The COVID-19 pandemic, seeing significant impacts from behaviour, also saw attempts to advance this field [15].
Despite this progress, important challenges remain in developing a generalised model of epidemiological and behavioural dynamics, including balancing model complexity, interpretability and capability to validate them with empirical data [16]. And at the same time, few models have attempted to capture specific behaviours, preferring to adopt an intervention-agnostic approach.
@@ 1.2 Outline of this study
Since testing and isolation are known to have both a strong behavioural component and a significant impact on disease transmission [17, 18], we define and analyse an extension to the Behaviour and Disease (BaD) model of Ryan et al. [19, 20] reflecting behavioural dynamics specific to (symptomatic) testing and isolation.
Specifically, we assume that the perception of illness threat reflects the rate of positive testing for disease, and thus we take the rate of behavioural uptake to be a function of that of positive testing, leading to different dynamics from previous models. In terms of data availability, it is also possible to monitor both test results and compliance with testing policy over time using community surveys [21], opening up possibilities for direct fitting of our model to data.
We structure the remainder of the paper as follows. We initially present our model that incorporates dynamic testing behaviour and infectious disease transmission processes (Section˜2). We then conduct theoretical analyses of the model (Section˜3) to give insight on the early dynamics and endemicity of the system; specifically, we derive its basic reproduction number $\mathcal{R}_{0}$ (Section˜3.1), analyse stability criteria (Section˜3.2) and compute its steady states (Section˜3.3). Note that derivation of quantities such as final size in the absence of waning immunity are not believed to be easily available for this class of models using existing methods (Röst, personal communication relating to the model published in Sayyar et al. [22]). Accordingly, to investigate how temporal and cumulative outbreak metrics depend on model parameter values we describe the results of numerical simulations (Section˜4). We conclude by discussing the findings and their implications (Section˜5).
While our analysis is intended to demonstrate overall features of our model, we note that most novelty arises from the modelled dependence of behavioural uptake due to perception of illness threat on positive testing rates. We will see that it is possible to simulate scenarios where increasing linear dependence of behavioural uptake on positive testing (captured by a parameter we call $\omega_{2}$) can lead to epidemics that have higher reported numbers of cases, but are objectively better mitigated in terms of having lower numbers of infections. This phenomenon is shown in Figure 4 (b,e,h).
@@ 2 Behaviour and Disease (BaD) model for symptomatic testing
In this section we introduce our mathematical model, which extends the BaD model of Ryan et al. [19] to incorporate dynamics of symptomatic testing. Ryan et al.’s BaD model considers a susceptible-infectious-recovered-susceptible infection process paired with a non-specific visual protective health behaviour. We extend this work in two ways. First, we consider more complex pathogen dynamics, including asymptomatic infection, as described below. This is primarily a step towards our main aim and second extension of previous work, which is to investigate the specific protective behaviour of symptomatic testing. Both generalisations are shown in the compartmental diagram for our model, Figure 1.
@@
A key concept behind our model is the distinction between individuals who, at any particular time, do not intend to seek testing when symptomatic, referred to as non-behavers (labelled $N$) and those who may seek testing when symptomatic, referred to as behavers (labelled $B$). In general, these intentions may change over time. We emphasise, however, that while we use the nomenclature of behavers and non-behavers for consistency with the other literature, we are considering $B$ individuals as those who hold an intention to test if symptomatic, and so may fail to do this if, for example, there is a lack of testing capacity. This means we parameterise the model in terms of the effectiveness of testing as an intervention, including the real-world factors that can obstruct access to testing, rather than its efficacy under ideal conditions that is primarily determined by test sensitivity.
@@
Our pathogen dynamical framework is built upon the standard SEIRS compartmental framework: susceptible individuals ($S$) become exposed ($E$) at a rate $\lambda$ (defined below), progress to an infectious state ($I$, AA or $T$ as explained below) at a rate $\sigma$ and then recover ($R$) at a rate $\gamma$, with a non-zero rate of waning immunity $\nu$ meaning that recovered individuals can become susceptible again. The rate $\lambda$ captures the force of infection, whereas $\sigma^{-1}$, $\gamma^{-1}$ and $\nu^{-1}$ capture the average latent, infectious, and immune periods, respectively. To account for symptomatic testing behaviour, we split the infectious class into asymptomatics (or pauci-symptomatics) labelled $A$ with probability $p_{A}$, and symptomatics. Symptomatics are further subdivided and are labelled $T$ with probability $p_{T}$ if they seek and receive a positive test, and $I$ otherwise. The value $p_{A}$ captures the probability of being asymptomatic, whereas $p_{T}$ captures the effectiveness of the test, as defined previously. The transitions between epidemiological states follow the standard SEIRS framework, while the changes in testing behaviour are governed by the rates of behavioural uptake ($\omega$) and abandonment ($\alpha$).
@@
Let $B(t)$ denote the proportion of the population at time $t$ that intend to test if showing symptoms, and $N(t)$ denote the proportion that do not. We will henceforth mostly suppress explicit dependence on $t$ for notational compactness. For each state class $X\in\{S,E,I,A,R\}$, $X_{N}$ and $X_{B}$ denote the proportions of non-behavers and behavers, respectively. Additionally, let the proportion of symptomatic individuals who seek testing and test positive be $T$.
@@
We adopt similar functional forms for the rates of behavioural uptake and behaviour loss as in the original BaD model [19], with one key difference in behavioural uptake. The rate of behavioural uptake is governed by
@@
where $\phi_{1}$ ($\omega_{1}$) represents the influence of social contagion on testing behaviour, $\phi_{2}$ ($\omega_{2}$) accounts for behavioural uptake driven by the perception of illness threat, and $\omega_{3}$ captures spontaneous adoption of the behaviour. Here, perception of illness threat is influenced by the prevalence of positive tests in the population ($T$, highlighted in Figure 1) as opposed to the prevalence of infection ($I_{N}+I_{B}+T$) in the BaD model [19], which also did not capture asymptomatics. The rate of behaviour loss is
@@
where $\psi$ ($\alpha_{1}$) represents the influence of social contagion on not seeking testing and $\alpha_{2}$ captures spontaneous abandonment of the behaviour. Note that while, for a complex contagion model, the quadratic and higher-order terms on the right-hand sides of Equations (1) and (2) might be large and even provide the dominant contribution to $\omega$ and $\alpha$, we will only include linear terms in our analysis when an explicit form is required.
@@
whereas for behavers it is $q_{B}\lambda$. Here, $\beta$ is the transmission rate for the pathogen while the values $q_{B},q_{A}$ and $q_{T}$ capture the relative reduction in infectiousness for behavers, asymptomatics, and test-and-isolators respectively.
Figure 1: A compartmental diagram illustrating the BaD model for symptomatic testing. We distinguish between individuals who do not seek testing when symptomatic, referred to as non-behavers (labelled $N$) and those who intend to seek testing when symptomatic, referred to as behavers (labelled $B$). Changes in testing behaviour are governed by the rates of behavioural uptake ($\omega$) and abandonment ($\alpha$). Uptake due to perception of illness threat is with respect to those in the highlighted compartment $T$. The epidemiological states are susceptible ($S$), exposed ($E$), asymptomatic infectious (or pauci-symptomatics) labelled AA, symptomatic infectious who are labelled II if they do not seek or do not receive a positive test and $T$ if they seek and receive a positive test, and those who are recovered ($R$). The transitions between epidemiological states follow a standard SEIRS framework, with waning immunity where recovered individuals can become susceptible again. We define the model parameters in Table 1.
@@
    $q_{A}$ the reduction in transmission due to being asymptomatic $0.58$ [24]
    $q_{T}$ the reduction in transmission due to (imperfect) isolation $0.25$
    $q_{B}$ the reduction in transmission due to other changes in behaviour from being willing to test $0.5$ [19]
@@
$\omega(t)$ the rate of testing behavioural uptake
@@
$\alpha(t)$ the rate of testing behaviour abandonment
@@
with B=1NB=1-N. Linearising about small initial behaviour ($B(0)\ll 1$) we have $N\approx 1$ and, if $\omega_{3}=0$, behaviour spreads through social pressure if $\omega_{1}>\alpha_{1}+\alpha_{2}$. Following Ryan et al. [19], we define the social reproduction number to be
@@
$F(N)=\left(\alpha_{1}-\omega_{1}\right)N^{2}-\left(\alpha_{1}-\alpha_{2}-\omega_{1}-\omega_{3}\right)N-\alpha_{2}=0\,.$ (14)
@@ 3 Theoretical analysis
In this section, we first derive the basic reproduction number $\mathcal{R}_{0}$ (Section˜3.1), then establish stability conditions for the infection-free equilibrium (Section˜3.2), and finally characterise the endemic steady states (Section˜3.3).
@@ 3.1.1 In the absence of behaviour
Consider when $\omega_{3}=0$ and $\mathcal{R}_{0}^{B}<1$. Then behaviour will be absent from the model and the infection-free steady state has $S_{N}=1$ and $S_{B}=0$. In this situation, the model simplifies to the standard SEIRS model with asymptomatics, and so the epidemic takes off if
@@ 3.1.2 In the presence of behaviour
We now consider the full model where both behaviour and infection are present (Equations (3)-(13)). Here we calculate the next-generation matrix, which will turn out to be equal to
$\mathbf{K}=\left(\begin{array}[]{cc}\Lambda_{N}N_{0}&\Lambda_{B}N_{0}\\ q_{B}\Lambda_{N}B_{0}&q_{B}\Lambda_{B}B_{0}\end{array}\right),$ (15)
where $B_{0}$ and $N_{0}$ are the proportions of the population behaving and not respectively at the start of the epidemic, $q_{b}$ is the reduction in transmission due to other changes in behaviour from being willing to test, and $\Lambda_{B}$ and $\Lambda_{N}$ are the forces of infection produced by behavers and non-behavers respectively. The dominant Eigenvalue of this matrix will give the basic reproduction number that we present and explain in Equation (23) below.
We now derive the result above by considering the different transmission and transition events possible. The model has two states-at-infection, $E_{N}$ and $E_{B}$. Let $t_{N}(X)$ denote the expected time an individual spends in compartment $X$ following state-at-infection $E_{N}$. Similarly, let $t_{B}(X)$ represent the corresponding time spent in compartment $X$ following state-at-infection $E_{B}$.
Defining $\alpha_{0}=\alpha(N_{0})$ and $\omega_{0}=\omega(B_{0},0)$, we derive expressions for the expected time spent in each compartment. Consider an individual with state-at-infection $E_{N}$ spending time in the asymptomatic state $A_{N}$. This individual will exit the exposed state after $\sigma^{-1}$ time units, on average, enter the asymptomatic state at a rate $p_{A}\sigma$, and will spend on average $\gamma^{-1}$ time units in the asymptomatic state. However, there are several possible paths this individual may take to move from $E_{N}$ to $A_{N}$. They may move directly from $E_{N}$ to $A_{N}$, or they may pass through $E_{B}$ (possibly alternating between $E_{B}$ and $E_{N}$ some number of times) to reach $A_{N}$. For an individual starting in $E_{N}$, the probability they exit $E_{N}$ into $A_{N}$ is
$\frac{\sigma}{\sigma+\omega_{0}}p_{A}\sum_{i=0}^{\infty}\left\{\frac{\alpha_{0}\omega_{0}}{(\sigma+\alpha_{0})(\sigma+\omega_{0})}\right\}^{i}=\frac{\sigma+\alpha_{0}}{\sigma+\omega_{0}+\alpha_{0}}p_{A}$
where the term $p_{A}\sigma/\left(\sigma+\omega_{0}\right)$ captures the probability of exiting epidemiologically out of $E_{N}$ into $A_{N}$, and the infinite sum accounts for any behavioural loops to $E_{B}$. Alternatively, the probability that they exit the exposed state via $E_{B}$ into $A_{B}$ is
$\left\{1-\frac{\sigma+\alpha_{0}}{\sigma+\omega_{0}+\alpha_{0}}\right\}p_{A}=\frac{\omega_{0}}{\sigma+\omega_{0}+\alpha_{0}}p_{A}\,.$
This individual will then transition from $A_{B}$ to $A_{N}$ with probability
$\frac{\alpha_{0}}{\alpha_{0}+\gamma}.$
Finally the individual transitions epidemiologically from $A_{N}$ to $R_{N}$ (again allowing for behavioural loops to $A_{B}$) with probability
$\frac{\gamma+\alpha_{0}}{\gamma+\omega_{0}+\alpha_{0}}\,.$
Thus, combining the behavioural and epidemiological aspects and simplifying, we find
$t_{N}(A_{N})$ $=\frac{p_{A}\left[\alpha_{0}\left(\sigma+\alpha_{0}+\omega_{0}\right)+\gamma\left(\sigma+\alpha_{0}\right)\right]}{\gamma\left(\sigma+\alpha_{0}+\omega_{0}\right)\left(\gamma+\alpha_{0}+\omega_{0}\right)}\,.$ (16)
Applying the same process to all other infectious states, for individuals with state-at-infection $E_{N}$ we have
$t_{N}(I_{N})$ $=\frac{\left(1-p_{A}\right)\left[\alpha_{0}\left(\sigma+\alpha_{0}+\omega_{0}\right)+\gamma\left(\sigma+\alpha_{0}\right)-p_{T}\alpha_{0}\omega_{0}\right]}{\gamma\left(\sigma+\alpha_{0}+\omega_{0}\right)\left(\gamma+\alpha_{0}+\omega_{0}\right)}$ (17)
@@
By substituting Equations (16)–(20), the terms $\Lambda_{N}$ and $\Lambda_{B}$ can be further expressed as
@@
where $\Lambda_{N},\Lambda_{B}$ are given by Equations (21) and (22), $N_{0}$ solves Equation (14) and $B_{0}=1-N_{0}.$ Note that $\Lambda_{N}$ and $\Lambda_{B}$ can be rewritten as
@@
where $\mathcal{R}_{0}^{D}$ is the behaviour-free reproduction number. Defining
$\Delta:=\frac{\beta(1-p_{A})(1-q_{T})p_{T}}{\gamma}$
enables us to express $\mathcal{R}_{0}$ in a form that makes explicit the reductions in transmission due to the different behavioural aspects of the model, such that
$\mathcal{R}_{0}=\underbrace{N_{0}\mathcal{R}_{0}^{D}}_{\text{(i)}}-\underbrace{(1-p_{N})N_{0}\Delta}_{\text{(ii)}}+\underbrace{q_{B}B_{0}\mathcal{R}_{0}^{D}}_{\text{(iii)}}-\underbrace{q_{B}p_{B}B_{0}\Delta}_{\text{(iv)}}\,.$ (23)
In Equation (23): (i) captures the spread of infection from non-isolators (those in $A$ and $I$) to non-behavers (those in $N$); (ii) captures the reduction in infections in non-behavers caused by isolation (those in $T$); (iii) captures the spread of infection from non-isolators to behavers (those in $B$), and; (iv) captures the reduction in infections in behavers caused by isolation. Note that in the limiting case where $B_{0}$ approaches 0, $p_{N}$ approaches $1$ and Equation (23) shows that $\mathcal{R}_{0}$ approaches $\mathcal{R}_{0}^{D}$.